OpenStudy (anonymous):

What is the equation of a line that is parallel to the line y=3/4x+2 and passes through the point (-2, 8)?

OpenStudy (anonymous):

Parallel --> same slope. slope here is 3 / 4 plug in point for x & y to solve for b (y-intercept) y = 3/4 x + b 8 = (3/4)(-2) +b 8 = (-6/4) + b 8 = (-3/2) + b 8 + 3/2 = b *side note 8 = 16/2 * 16/2 + 3/2 = b 19/2 = b 9.5 = b So equation is: y = 3/4 (x) + 9.5 Shhh... · 4 years ago

OpenStudy (anonymous):

IM pretty sure @acosta2943

OpenStudy (anonymous):

Thanks Hannah

OpenStudy (anonymous):

Welcome :) @acosta2943